OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

2. The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

let Tim is x years old johnson is then x+10 years old 2(x-3)=(x+10-3) 2x-6=x+7 x=13 therefore, Tim is 13 now

OpenStudy (anonymous):

let the largest integer be x then the other 2 numbers are x-1 and x-2 x+x-1+x-2=267 3x-3=267 3x=270 x=90 so the largest number is 90